### 26 Feb 2024 v4

# Diffusion Model with Perceptual Loss

Abstract
Diffusion models trained with mean squared error loss tend to generate unrealistic samples. Current state-of-the-art models rely on classifier-free guidance to improve sample quality, yet its surprising effectiveness is not fully understood. In this paper, we show that the effectiveness of classifier-free guidance partly originates from it being a form of implicit perceptual guidance. As a result, we can directly incorporate perceptual loss in diffusion training to improve sample quality. Since the score matching objective used in diffusion training strongly resembles the denoising autoencoder objective used in unsupervised training of perceptual networks, the diffusion model itself is a perceptual network and can be used to generate meaningful perceptual loss. We propose a novel self-perceptual objective that results in diffusion models capable of generating more realistic samples. For conditional generation, our method only improves sample quality without entanglement with the conditional input and therefore does not sacrifice sample diversity. Our method can also improve sample quality for unconditional generation, which was not possible with classifier-free guidance before.
footnotetext: Model: hf.co/ByteDance/sd2.1-base-zsnr-laionaes6-perceptualfootnotetext: Code: See Algorithm 1.
1 Introduction
Diffusion models [10, 35, 37] are a rising class of generative model. Conceptually, they work by converting pure noise to a data sample through repeated denoising. Formally, each denoising step can be viewed from the lens of score matching [37] such that the model learns to predict the gradient (score) of an ordinary differential equation (ODE) or a stochastic differential equation (SDE) that transforms samples from one distribution (noise) to another distribution (image, video, etc.) [18]. In this paper, we focus on image generation but the findings are also applicable to other modalities.
Diffusion models are commonly parameterized as neural networks and the training objective minimizes the squared distance between the model prediction and the ground truth score through stochastic gradient descent [12]. This is also commonly referred to as the mean squared error (MSE) loss.
Although diffusion models are supposed to transport samples from noise to image distribution by theory, images generated by a diffusion model in its raw form are often of poor quality, despite the improvements in model architecture [4, 14, 22, 24, 29, 27, 28], formulation [17, 18, 13], and sampling strategy [36, 19, 20, 13].
What drives diffusion models into the mainstream is the advent of classifier guidance [5] and classifier-free guidance [11]. Classifier guidance shows that we can add classifier gradients on top of the predicted score during the inference process to guide the sample generation toward the classifier’s direction. Therefore, it can turn an unconditional diffusion model conditional. However, it is surprising that applying classifier guidance on an already conditional diffusion model can significantly improve sample quality. Classifier-free guidance [11] improves classifier guidance by removing the need for an external classifier network. It simultaneously trains the diffusion model as both conditional and unconditional by dropping the condition with a certain probability. At inference, it queries the model both conditionally and unconditionally at every step and uses their difference as the conditional gradient direction. The score prediction is then amplified toward this conditional direction.
Classifier-free guidance is applied almost ubiquitously in state-of-the-art diffusion models across modalities, e.g. text-to-image [24, 28, 29, 4, 27, 45, 23], text-to-video [7, 2, 46, 9, 34, 1], text-to-3d [33, 41, 25], image-to-video [1, 43], video-to-video [3, 6], etc. Yet, its surprising effectiveness in improving sample quality is not fully understood. Classifier-free guidance also has many limitations, such as having a sensitive hyperparameter that can cause image over-exposure and over-saturation, etc. [15, 29], but there are no other viable alternatives.
In this paper, we elucidate that the effectiveness of classifier-free guidance in improving sample quality partly originates from it being an implicit form of perceptual guidance. We show that perceptual loss can be directly applied to diffusion training to improve sample quality. Instead of using external perceptual networks, we find the diffusion model itself is already a well-trained perceptual network. We propose a novel self-perceptual objective that utilizes the diffusion model itself to generate meaningful perceptual loss. Our training objective results in an improvement in sample quality. Unlike classifier-free guidance, our method does not rely on conditional input and therefore does not trade off sample diversity.
2 Problem
2.1 Diffusion Models Generate Bad Samples
Let’s first analyze why diffusion models in their raw form generate bad samples. As illustrated in Figure 0(a), consider a simple distribution where the images always contain a solid circle at arbitrary locations. Given the model only a few training samples (top row), we want the model to generate a new sample from this distribution (bottom left), but the actual generation can be far out of the distribution (bottom right).
From a non-biased standpoint, the model’s generation of two circles is equally plausible given the limited training samples. When observations are sparse in a distribution or a sub-region of a distribution, there is room for ambiguities.
Another factor that can cause the issue is model capacity. Even when more training samples are given, the model must have enough capacity to learn and deduce that the underlying distribution most likely always contains one circle only.
Figure 1: Illustration of the current diffusion model problem. Diffusion models learn a plausible pixel distribution from the training samples but it does not align with the distribution of real images.
As illustrated in Figure 0(b), similar problems exist for actual images. The diffusion model tries to learn a plausible pixel distribution from finite training samples, but it does not always align with the intended distribution of photorealistic images.
Existing works have focused on expanding the dataset [32, 4], increasing the model capacity [24, 4, 14, 22, 29, 27], simplifying the ODE trajectory [18, 17, 13], and improving the sampler [36, 19, 20]. However, almost all works use the mean squared error (MSE) loss in training and no study has explored the use of perceptual loss.
2.2 Perceptual Loss is the Hidden Gem
Prior works have shown that the mean squared error (MSE) metric aligns poorly with human perception [40, 42, 44]. For example, when comparing two faces, humans are much more sensitive if there is a mismatch between the eyes than the hair strands. As another example, shifting an image by a few pixels is almost undetectable by human perception but can cause a large MSE value. Thus, when training diffusion models with MSE loss, it penalizes the imperceivable pixel mismatch more than perceivable structural features. This is clearly not an ideal use of model capacity.
Prior works have found that the distance computed on the hidden features of deep neural networks can be used as a metric that resembles human perception better than the distance computed directly on image pixels [44]. This is because deep neural networks can learn high-level semantics rather than only focusing on pixel differences.
In fact, we will show that this neural perceptual property has already been implicitly applied in classifier guidance [5] and classifier-free guidance [11] in the following sections.
2.3 Perception in Classifier Guidance
Originally, classifier guidance [5] was proposed to guide an unconditional diffusion model to generate class conditional samples, but prior works have found that it can also help an already conditional diffusion model to improve sample quality and attributed it as a side effect of trading off sample diversity.
However, we argue that the surprising effectiveness of classifier guidance also partially originates from it being an implicit form of perceptual guidance. Specifically, the classifier network, i.e. CLIP [26], is a trained semantic neural network that can provide perceptual guidance. The guidance gradients are high on more perceptually important areas, such as the shape and contour, and low on the imperceivable areas. The diffusion model is implicitly guided toward what the perceptual classifier considers as more probable, resulting in more perceptually photorealistic samples.
2.4 Perception in Classifier-Free Guidance
Classifier-free guidance (CFG) [11] can also be viewed from the lens of perceptual guidance. The diffusion model itself is a neural network that learns high-level semantic features, and classifier-free guidance exploits the perceptual knowledge from the model itself.
Classifier-free guidance is defined as:
$x_{cfg}=x_{\oslash}+w(x_{c}-x_{\oslash})$ (1)
We first query the model unconditionally $x_{\oslash}$, then query the model conditionally $x_{c}$. We subtract them to get the “difference”, then accentuate this difference by a large factor $w$.
This “difference” is in fact the high-level semantic perception that is the most essential about the target prompt. The difference is large in more perceptually important areas and small in unimportant areas. In this implicit way, classifier-free guidance is a form of perceptual guidance as well.
2.5 Limitations of Classifier-Free Guidance
Classifier-free guidance has numerous limitations. This incentivizes us to explore directly incorporating perceptual loss in diffusion training. Specifically:
  • Classifier-free guidance only works for conditional models. Perceptual loss works for unconditional model as well.
  • Classifier-free guidance tangles text alignment with sample quality. Perceptual loss only improves sample quality and is condition agnostic.
  • Classifier-free guidance is added post training. High scale can lead to overexposure [15, 29]. Perceptual loss is applied at training and does not suffer this issue.
3 Method
Our goal is to incorporate perceptual loss into diffusion training to improve sample quality. In Section 3.1 we introduce the diffusion background and our model formulation. In Section 3.2, we propose a novel self-perceptual objective and show that the diffusion model itself can be used as a perceptual network to provide meaningful perceptual loss.
3.1 Background
We follow the setup of Stable Diffusion, a latent diffusion model [28]. Given image latent samples of $x_{0}\sim\pi_{0}$, noise samples of $\epsilon\sim\mathcal{N}(0,\textbf{I})$, and time $t\sim\mathcal{U}(1,T)$, where $t\in\mathbb{Z},T=1000$, the forward diffusion process is defined as:
$x_{t}=forward(x_{0},\epsilon,t)=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$ (2)
We use diffusion schedule with zero terminal SNR fix [15]. The specific $\bar{\alpha}_{t}$ values are defined in [15].
Our neural network $f_{\theta}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is conditioned on text prompt $c$ and uses the $v$-prediction formulation [31, 15]:
$v_{t}=\sqrt{\bar{\alpha}_{t}}\epsilon-\sqrt{1-\bar{\alpha}_{t}}x_{0}$ (3)
$\hat{v}_{t}=f_{\theta}(x_{t},t,c)$ (4)
With vanilla diffusion training objective, we optimize the following MSE loss:
$\mathcal{L}_{mse}=\|\hat{v}_{t}-v_{t}\|_{2}^{2}$ (5)
3.2 Self-Perceptual Objective
Prior works have shown that the score matching objective strongly resembles the denoising autoencoder objective [38]. Denoising autoencoder objective is often used for unsupervised pre-training of neural networks [39]. Thus the diffusion model trained with vanilla MSE loss is effectively a perfect unsupervised perceptual network trained on the target dataset, on the latent space, and on all noise levels $x_{t}$. In this section, we show that we can exploit the diffusion model itself as a perceptual network to provide meaningful perceptual loss.
First, we copy and freeze the diffusion model trained with vanilla MSE loss, and modify the architecture to return hidden feature at layer $l$. We denote this frozen network as $f^{l}_{*}$.
During training, we sample $x_{0}\sim\pi_{0}$, $\epsilon\sim\mathcal{N}(0,\textbf{I})$, $t\sim\mathcal{U}(1,T)$ and compute $x_{t}$ through forward diffusion:
$x_{t}=forward(x_{0},\epsilon,t)$ (6)
We use online network $f_{\theta}$ to predict $\hat{v}$ and convert the prediction to $\hat{x}_{0}$ and $\hat{\epsilon}$:
$\hat{v}_{t}=f_{\theta}(x_{t},t,c)$ (7)
$\hat{x}_{0}=\sqrt{\bar{\alpha}_{t}}x_{t}-\sqrt{1-\bar{\alpha}_{t}}\hat{v}$ (8)
$\hat{\epsilon}=\sqrt{\bar{\alpha}_{t}}\hat{v}+\sqrt{1-\bar{\alpha}_{t}}x_{t}$ (9)
Then, we sample a new $t^{\prime}\sim\mathcal{U}(1,T)$, and compute $x_{t^{\prime}}$ and $\hat{x}_{t^{\prime}}$ through forward diffusion:
$x_{t^{\prime}}=forward(x_{0},\epsilon,t^{\prime})$ (10)
$\hat{x}_{t^{\prime}}=forward(\hat{x}_{0},\hat{\epsilon},t^{\prime})$ (11)
Finally, we pass them through the frozen network $f^{l}_{*}$ and compute MSE on its hidden feature at layer $l$. We find only using the hidden feature at midblock layer yields the best result. We refer to our method as the Self-Perceptual (SP) objective:
$\mathcal{L}_{sp}=\|f^{l}_{*}(\hat{x}_{t^{\prime}},t^{\prime},c)-f^{l}_{*}(x_{t^{\prime}},t^{\prime},c)\|_{2}^{2}$ (12)
The pseudo code is provided in Algorithm 1.
4 Evaluation
We first finetune Stable Diffusion v2.1 [28] using our formulation and MSE loss $\mathcal{L}_{mse}$ on a subset of LAION dataset [32] where the images have resolution greater than 512px and aesthetic score above 6. We use learning rate 3e-5, batch size 896, ema decay 0.9995 for 60k iterations. We also use 10% conditional dropout to support CFG for evaluation comparison. Then we copy and freeze the model as our perceptual network, and continue training the online network with our self-perceptual objective $\mathcal{L}_{sp}$ for 50k iterations. This does not have a conditional dropout.
For inference, we use deterministic DDIM sampler [36], and make sure the sampler starts from the last timestep $T$ [15].
4.1 Qualitative
Figure 2 shows example generation results. Our self-perceptual objective has visible quality improvement over the vanilla MSE objective, but the overall sample quality is still worse than classifier-free guidance. This is because our objective only improves sample quality whereas classifier-free guidance has the additional effect of enhancing text alignment. This is especially obvious in Figure 1(j), where our objective only enhances the sample quality without extra emphasis on the text condition.
Notice that the results of vanilla MSE and self-perceptual objective share similar image content and layout when generated from the same initial noise, whereas classifier-free guidance will largely alter the result. Our self-perceptual objective maintains the same diversity whereas classifier-free guidance does not.
Additionally, Figure 1(i) shows the negative artifact of classifier-free guidance. The model has already overfitted the image to the very specific prompt, and the high classifier-free guidance scale causes unnatural artifacts. Our self-perceptual objective does not suffer from this issue.
MSE
Self-Perceptual
MSE + CFG
MSE
Self-Perceptual
MSE + CFG
(a) New York Skyline with ‘Deep Learning’ written with fireworks on the sky.
(b) An elephant is behind a tree. You can see the trunk on one side and the back legs on the other.
(c) A couple of glasses are sitting on a table.
(d) One car on the street.
(e) A laptop on top of a teddy bear.
(f) A giraffe underneath a microwave.
(g) A yellow book and a red vase.
(h) One cat and one dog sitting on the grass.
(i) A painting by Grant Wood of an astronaut couple, american gothic style.
(j) A blue colored dog.
(k) McDonalds Church.
(l) Photo of a cat singing in a barbershop quartet.
Figure 2: Text-to-image generation on DrawBench prompts [29]. Our self-perceptual objective improves sample quality over the vanilla MSE objective while largely maintaining the image content and layout. Classifier-free guidance has the additional effect of enhancing text alignment by sacrificing sample diversity. Images are generated with DDIM 50 NFEs. More analysis in Sec. 4.1.
MSE
Self-Perceptual
MSE + CFG
MSE
Self-Perceptual
MSE + CFG
(a) A black colored car.
(b) A black colored sandwich.
(c) A fish eating a pelican.
(d) A green apple and a black backpack.
(e) A green cup and a blue cell phone.
(f) A horse riding an astronaut.
(g) A pink colored giraffe.
(h) A red colored car.
(i) A separate seat for one person, typically with a back and four legs.
(j) A sign that says ‘Diffusion’.
(k) A sheep to the right of a wine glass.
(l) A photo of a confused grizzly bear in calculus class.
Figure 3: Text-to-image generation on DrawBench prompts [29]. Our self-perceptual objective improves sample quality over the vanilla MSE objective while largely maintaining the image content and layout. Classifier-free guidance has the additional effect of enhancing text alignment by sacrificing sample diversity. Images are generated with DDIM 50 NFEs. More analysis in Sec. 4.1.
4.2 Quantitative
Table 1 shows our quantitative evaluation. We follow the convention to calculate Fréchet Inception Distance (FID) [8, 21] and Inception Score (IS) [30]. We select the first 10k samples from the MSCOCO 2014 validation dataset [16] and use our models to generate images of the corresponding captions. Our self-perceptual objective has improved FID/IS over the vanilla MSE objective, this aligns with our observed improvement in sample quality. However, classifier-free guidance still achieves better sample quality than our self-perceptual objective.
Loss
CFG Rescale Steps NFE FID $\downarrow$ IS $\uparrow$
Ground truth
00.00 35.28
$\mathcal{L}_{mse}$
25 25 32.68 22.20
50 50 29.63 22.86
$\mathcal{L}_{mse}$
7.5 25 50 24.41 32.10
7.5 0.7 25 50 18.67 34.17
$\mathcal{L}_{sp}$
25 25 25.89 27.76
50 50 24.42 28.07
Table 1: Quantitative evaluation on MSCOCO 10K validation dataset. Our self-perceptual (SP) objective improves FID and IS metrics over vanilla MSE objective, but is still worse than using classifier-free guidance [11] with rescale [15]. This is expected because classifier-free guidance additionally enhances text alignment and trades off sample diversity. Since classifier-free guidance with 25 steps incurs 50 NFEs (number of function evaluations), we show both 25 steps and 50 steps metrics.
5 Ablation Study
In this section, we evaluate the individual hyperparameters we choose for our self-perceptual objective. All metrics are calculated on the same MSCOCO 10k validation samples as in Section 4.2 and use 25 steps DDIM inference.
5.1 Layer $l$
Layer
FID $\downarrow$ IS $\uparrow$
All Encoder Layers
26.64 26.89
All Decoder Layers
42.42 19.98
All Encoder Layers + Midblock Layer
26.96 27.24
Only Midblock Layer
25.89 27.76
Table 2: Comparing computing perceptual loss on different layers. We find that only computing loss on the midblock hidden features yields the best results.
We compare the effect of computing loss on hidden features from different layers $l$. We try using all the hidden features from every encoder layer and decoder layer by summing up the losses. We also try using only features from the midblock layer. As shown in Table 2, we find that only using features from the midblock layer yields the best metrics.
5.2 Timestep $t^{\prime}$
We compare the effect of choosing timestep $t^{\prime}$ for the perceptual network. First, we show that $t^{\prime}=t$ is invalid because $\hat{x}_{t}$ always equals $x_{t}$. This will make the input to the perceptual network identical which prevents any meaningful loss:
$\hat{x}_{t}=forward(\hat{x}_{0},\hat{\epsilon},t)=forward(x_{0},\epsilon,t)=x_{t}$ (13)
We compare three different choices for $t^{\prime}$. Table 3 shows that uniform sampling of $t^{\prime}$ yields the best results.
Timestep ($t^{\prime}$ clamped to $[1,T]$)
FID $\downarrow$ IS $\uparrow$
$t^{\prime}=t\pm 40$
27.24 23.31
$t^{\prime}\sim\mathcal{N}(t,100)$
24.54 25.42
$t^{\prime}\sim\mathcal{U}(1,T)$
25.89 27.76
Table 3: Comparing the choice of timestep $t^{\prime}$. We find that simply uniformly sample $t^{\prime}$ can yield reasonably good results.
5.3 Distance Function
We compare using different distance functions on the hidden features. Table 4 shows that MSE and MAE yield very similar results, so we stick to MSE.
Distance FID $\downarrow$ IS $\uparrow$
Mean Absolute Distance
($\|\cdot\|_{1}$)
25.28 27.41
Mean Squared Distance
($\|\cdot\|_{2}^{2}$)
25.89 27.76
Table 4: Comparing the choice of distance function. We find that mean squared distance and mean absolute distance have similar results, so we stick to mean squared distance.
5.4 Formulation
We experiment with an alternative formulation, which combines predicted $\hat{x}_{t^{\prime}},\hat{\epsilon}_{t^{\prime}}$ separately with ground-truth $x_{t^{\prime}},\epsilon_{t^{\prime}}$. This formulation allows gradient feedback at $t^{\prime}=t$:
$x_{t^{\prime}}=forward(x_{0},\epsilon,t^{\prime})$ (14)
$\hat{x}_{t^{\prime}}^{x}=forward(\hat{x}_{0},\epsilon,t^{\prime})$ (15)
$\hat{x}_{t^{\prime}}^{\epsilon}=forward(x_{0},\hat{\epsilon},t^{\prime})$ (16)
$\mathcal{L}_{sp2}=$ $\|f^{l}_{*}(\hat{x}_{t^{\prime}}^{x},t^{\prime},c)-f^{l}_{*}(x_{t^{\prime}},t^{\prime},c)\|_{2}^{2}$ (17)
$+$ $\|f^{l}_{*}(\hat{x}_{t^{\prime}}^{\epsilon},t^{\prime},c)-f^{l}_{*}(x_{t^{\prime}},t^{\prime},c)\|_{2}^{2}$
Table 5 shows that the alternative formulation yields worse performance.
Formulation
FID $\downarrow$ IS $\uparrow$
$\mathcal{L}_{sp}$
25.89 27.76
$\mathcal{L}_{sp2}$
29.83 24.54
Table 5: Comparing different formulation. We find that merged formulation yields the best results.
5.5 Repeat Perceptual Network
We experiment using the network trained with self-perceptual objective as the perceptual metric network $f^{l}_{*}$ and repeat the training process. Table 6 shows that repeating the self-perceptual training results in worse performance. This is why we decide to just freeze the MSE model instead of using an exponential moving average (EMA) for the perceptual network.
Formulation
FID $\downarrow$ IS $\uparrow$
MSE model as perceptual network
25.89 27.76
SP model as perceptual network
26.61 26.41
Table 6: Repeating the self-perceptual process yields worse performance.
5.6 Combine with Classifier-Free Guidance
We experiment with applying classifier-free guidance on the model trained with our self-perceptual objective. Table 7 shows that classifier-free guidance indeed can improve sample quality further on the self-perceptual model but it does not surpass classifier-free guidance applied on the MSE model.
Loss
CFG Rescale FID $\downarrow$ IS $\uparrow$
$\mathcal{L}_{mse}$
7.5 0.7 18.67 34.17
$\mathcal{L}_{sp}$
25.89 27.76
2.0 0.7 21.19 32.22
3.0 0.7 20.65 33.49
4.0 0.7 20.67 33.34
7.5 0.7 23.49 31.64
Table 7: Combining our self-perceptual objective with classifier-free guidance does improve sample quality but does not surpass the MSE objective with classifier-free guidance.
5.7 Unconditional Generation
MSE
Self-Perceptual
Figure 4: Unconditional generation. Both use DDIM 1000 steps with the same seed. Our self-perceptual objective can improve unconditional generation quality. This was previously not possible with classifier-free guidance because it only works for conditional models. More analysis in Sec. 5.7.
1000
900
800
700
600
500
400
300
200
100
Final
(a) MSE
(b) MSE + CFG
(c) Self-Perceptual
Figure 5: Model prediction at each step during inference converted to $\hat{x}_{0}$ space. This shows that the model trained with perceptual loss behaves very differently than MSE loss. More analysis in Sec. 5.8.
We train an unconditional image generation model following the same procedure except we always use an empty prompt during training and inference.
Figure 4 shows that the MSE objective generates unrealistic images even when using 1000 sample steps, whereas the self-perceptual objective generates much preferable results. This validates that our self-perceptual objective can improve the quality for unconditional diffusion models, which was not possible with classifier-free guidance before.
Table 8 shows the quantitative metric, which also shows that the self-perceptual objective is effective in improving unconditional generation quality.
Loss
IS $\uparrow$
$\mathcal{L}_{mse}$
11.18
$\mathcal{L}_{sp}$
12.04
Table 8: Unconditional generation metric. Self-perceptural objective improves inception score.
5.8 Inference Behavior
In Figure 5, we visualize the model prediction by converting to $\hat{x}_{0}$ at every inference step. We find that the model trained with self-perceptual generates shapes and contours earlier in the inference process. We also notice that it has grid-like pattern artifacts, likely resulting from the convolution downsampling nature of the perceptual network. This artifact does not affect overall generation much. We leave it for future investigation and improvement.
6 Conclusion
In summary, we have shown that the effectiveness of classifier guidance and classifier-free guidance can be viewed from the lens of perceptual guidance. We have found that perceptual loss can be directly applied to diffusion training to improve sample quality. Specifically, we have proposed a novel self-perceptual objective, which uses the diffusion model itself as a perceptual network. Our objective can be generalized to all modalities, i.e. image, video, audio, etc., and supports unconditional generation, which was not possible with classifier-free guidance before. For conditional generation, our objective improves sample quality without entanglement with conditional input.
However, for text-to-image generation, classifier-free guidance still generates overall better images than our self-perceptual objective. This is because classifier-free guidance has the additional effect of increasing text alignment by trading off diversity.
We hope our work paves the way for more future explorations on diffusion training loss.
1# Create dataloader
2dataloader = create_dataloader()
3
4# Create model by loading from mse pretrained weights.
5model = create_model(mse_pretrained=True)
6optimizer = Adam(model.parameters(), lr=3e-5)
7
8# Create perceptual model and freeze it.
9perceptual_model = deepcopy(model)
10perceptual_model.requires_grad_(False)
11perceptual_model.eval()
12
13# Dataloader yields image (latent) x_0, and conditional prompt c.
14for x_0, c in dataloader:
15
16    # Sample timesteps and epsilon noises.
17    # Then perform forward diffusion.
18    t = randint(0, 1000, size=[batch_size])
19    eps = randn_like(x_0)
20    x_t = forward(x_0, eps, t) # equation 2.
21
22    # Pass through model to get v prediction.
23    # Then convert v_pred to x_0_pred and eps_pred.
24    v_pred = model(x_t, t, c)
25    x_0_pred = to_x_0(v_pred, x_t, t) # equation 8.
26    eps_pred = to_eps(v_pred, x_t, t) # equation 9.
27
28    # Sample new timesteps.
29    # Then perform forward diffusion twice.
30    # One uses ground truth x_0 and eps.
31    # Another uses predicted x_0_pred and eps_pred.
32    tt = randint(0, 1000, size=[batch_size])
33    x_tt = forward(x_0, eps, tt)
34    x_tt_pred = forward(x_0_pred, eps_pred, tt)
35
36    # Pass through perceptual model.
37    # Get hidden feature from midblock.
38    feature_real = perceptual_model(x_tt, tt, c, return_feature="midblock")
39    feature_pred = perceptual_model(x_tt_pred, tt, c, return_feature="midblock")
40
41    # Compute loss on hidden features.
42    loss = mse_loss(feature_pred, feature_real)
43    loss.backward()
44    optimizer.step()
45    optimizer.zero_grad()
Algorithm 1 Pseudo PyTorch code for self-perceptual training.